I am the parent of a 5th grade student learning Algebra, I can solve the equations myself, and have no problem giving the answer, but how do I show my child how to show the work? They aren't teaching the same as they did when I was a student. Ex: 2x + 8 = 20 I know that x = 6, but how do I show how to get to the answer?
Answer by checkley79(3056)

2x + 8 = 20 subtract 8 from both sides.
2x=20-8 combine the right hand terms.
2x=12 divide both sides by 2.
x=12/2
x=6 ans.